OpenStudy (anonymous):

–5(1 – 5x) + 5(–8x – 2) = –4x – 8x A.x = –5 B.x = 0.19 C.x = 0.55 D.x = 5

OpenStudy (anonymous):

simplify it and then solve for x

OpenStudy (jgirl128):

did you simplify it yet?

OpenStudy (jgirl128):

–5(1 – 5x) + 5(–8x – 2) = –4x – 8x simplify -5 - 25x + -40x - 10 = -4x - 8x combine like terms -15x - 15 = -12x add 15x to both sides -15x - 15 = -12x +15x +15x ______ ______ -15 = 3x divide both sides by 3 -15 = 3x ____ ____ 3 3 now what would the answer be?

OpenStudy (jgirl128):

did you get the answer?

OpenStudy (anonymous):

yeah

OpenStudy (jgirl128):

so whats the answer?

OpenStudy (jgirl128):

I did all the work for you, I just wanna make sure you understand. all you have to do is divide -15 by 3.